# Visual Whole-Body Control for Legged Loco-Manipulation

### v4

@@ C.1. Privileged Teacher Policy
C.2. Visuomotor Student Policy We find robust and adaptive picking-up behaviors emerge from the privileged teacher policy training, with inaccessible object poses and shape features in reality. To deploy our policy in the real world, we must use easily accessible real-world observations like images.
Observations and actions. Although camera images are easily accessible both in simulation and the real world, there is a huge gap for colored images between the simulation and the real world. This persuades us to use depth images only. To further decrease the noise and gap from backgrounds, we use the segmented depth images and also their segmentation masks. Formally, the observation of our vision policy is
$\mathbf{o}_{t}=[\mathbf{o}^{\text{image}}_{t},\mathbf{s}_{t}^{\text{proprio}},\mathbf{a}_{t-1}],$
@@
where $\mathbf{o}^{\text{image}}_{t}$ consists of object segmentation masks and segmented depth images. $\mathbf{s}_{\text{proprio}}$ and $\mathbf{a}_{t-1}]$ are proprioceptive observation and last action defined above. The output actions are the same as those of state-based policy.
@@
Imitation learning for student distillation. To deploy our policy in the real world, we distill our state-based high-level planning policy into a vision-based policy using DAgger. We warm up by teacher sampling the initial transition data. This is like behavior cloning (BC) to find a good initialization of the visuomotor student policy to improve the training efficiency. Then we sample transitions using the student policy and request correction from the teacher. We find this is much more efficient and effective than directly learning a visuomotor student policy using RL. We adopt augmentations like RandomErasing, GaussianBlur, GaussianNoise, and RandomRotation for the depth images. More details can be found in Appendix II-B.
Policy network architecture. Our vision-based high-level policy network comprises a CNN and a two-layer MLP that outputs the high-level actions. We concatenate the images and other states as observations. We stack 4-step images which are passed through the two-layer CNN network with kernel size being $5$ and $3$, respectively, and then encoded as a $64$-dim latent vector. Then, we concatenate this latent vector with other states as a flattened vector. This vector is then passed to a two-layer MLP with a hidden size of 64. We use ELU as the activation.
III-D Real World Deployment
Figure 7: Segmented examples of the TrackingSAM tool. At the beginning of each reset trial, we annotated the objects from the two RGBD cameras mounted on the gripper and the robot head. Then, TrackingSAM will keep tracking the segmented object and providing the mask.
Hardware setup. As visualized in Fig. 4, we mount two cameras on our robot platform, which are both RealSense D435 cameras. B1 itself has an onboard computer, and we also attached an additional onboard computer, Nvidia Jetson Orin, to the robot. The B1, Z1, and the Orin computer are all powered by B1’s onboard battery. The inference of the low-level policy and the control of the arm is conducted at the onboard computer of B1, and that of the high-level vision-based planning policy is made on Nvidia Jetson Orin. It is worth noting that the computation and the power of our robot system are fully onboard and, therefore, untethered.
System design. During deployment, the low-level policy that controls the quadruped works at a frequency of 50Hz; in the meantime, the high-level policy runs at 8Hz, which is five times slower than the low-level policy; besides, the arm is required to work at a frequency of more than 800Hz. Since these policies work on different computers, along with the quadruped and the arm should work under different frequencies, we design a client-server multi-processing system for this work, as shown in Fig. 6. We set up three processes at the low-level policy side: a quadruped robot control process, an arm control process, and a goal receiver process behaving as a client. The high-level policy combines the vision inputs from the depth cameras and the information from the low-level request as the observation works as a server. Every 0.1 seconds, the goal receiver sends a request with the necessary information to the server; the high-level policy makes the latest inference and sends the high-level action back to the robot; at the same time, the quadruped and the arm control processes gather the latest high-level commands to execute, while keep updating the necessary information for the goal receiver process.
@@ III-D Real World Deployment
Sim-to-real transfer. We made several efforts to bridge the sim-and-real gap. Specifically, we slightly randomize the position and rotation of both our cameras to enable our policy to handle slight camera offsets from simulation. Besides, we clip the depth image to a minimum of 0.2m to avoid low-quality images when the distance is too close to a real camera, after which we normalize the depth value. As the surrounding environment in the simulation and the real world are very different, to reduce the noise, we only use segmented depth images to train our visuomotor policy.
@@
Vision masks. Note that our system works with only limited human intervention, i.e., we need to annotate and segment objects from environment backgrounds. To this end, we develop TrackingSAM [56], a tool that combines the tracking model AOT (Associating Objects with Transformers) [76] and SAM (Segment Anything) [34]. This allows us to annotate the object on the image from both cameras or an arbitrary image before starting the task and track that object to provide its real-time segmented masks. The tool is integrated into the high-level server, and a demo example can be found in Fig. 7. During real-world experiments, we annotated the objects from the two RGBD cameras mounted on the gripper and the robot head at the beginning of each reset trial. TrackingSAM will keep tracking the segmented object and providing the mask. The depth images obtained from the RGBD cameras are pre-processed by hole-filling and clipping. In real-world experiments, as long as the object is visible to one camera (either the head or the wrist camera), the annotation mask can be propagated to both views to initialize an automated pickup.
IV Experiments
We conduct a set of experiments on pick-up tasks to compare VBC with baselines both in simulation and the real world, showing effectiveness from the following perspectives:
  • VBC solves the mobile manipulation (pickup) task over varying objects with different heights;
  • VBC has the advantage over easy baseline methods, such as decoupling mobile manipulation problems into separate moving and arm-control;
  • The object shape feature produced by pre-trained PointNet++ boosts learned policy grasp varying objects;
Experiment setups.
@@ IV Experiments
The simulation / real-world setup is shown in Fig. 8. Particularly, in the simulation, the high-level policy is trained to pick up various objects on a range of table heights $[0,0.65]$ meter. We utilize IsaacGym [47] for parallel training in a large set of environments. In the real world, we initialize objects in a random location and orientation, and put them in front of the robot where it is visible through both cameras (around 0.5$\sim$1.5m). We test three different heights by putting the objects on the ground, on a box, and on a table, corresponding to the heights of 0.0m, around  0.3m,  0.5m.
Figure 9: Examples of training objects and real-world objects.
Figure 10: Success rates of different student policies at different heights, tested in the simulator. Each distinct object for more than 100 episodes. The robot starts from a 1.0-meter distance from the table. Every episode ends when the robot successfully picks up the object, the object falls from the table, or the robot fails until 150 steps ($\sim$ 18 seconds). The dots represent the mean performance of all objects, and the upper and the lower bound represent the maximum and the minimum success rate overall 7 categories, where we also note their category name.
IV-A Simulation Results
Objects. The objects used for the benchmark problems are shown in Fig. 9. Specifically, we included $34$ objects in total, mainly adapted from the YCB dataset [11]. According to the object shapes, we roughly divided them into $7$ categories: ball, long box, square box, bottle, cup, bowl and drill.
TABLE I: Success rates of VBC compared to baseline method on seven categories of objects, tested in the simulator. Each distinct object for more than 300 episodes. The robot starts from a 1.0-meter distance from the table. Every episode ends when the robot successfully picks up the object, the object falls from the table, or the robot fails until 150 steps. With the 3D shape feature, the algorithm is better at handling irregular objects. Detailed descriptions of each method are stated in Section IV-A.
@@ IV-A Simulation Results
Privleged Teacher Floating Base $63.25\%$ $20.51\%$ $72.66\%$ $38.07\%$ $69.55\%$ $71.88\%$ $4.17\%$
VBC w.o. Shape Feature $\mathbf{96.18}\%$ $74.26\%$ $\mathbf{96.80}\%$ $\mathbf{82.46}\%$ $84.38\%$ $55.84\%$ $73.77\%$
VBC (Ours) $89.08\%$ $\mathbf{92.83\%}$ $81.87\%$ $76.34\%$ $\mathbf{89.12\%}$ $\mathbf{78.14\%}$ $\mathbf{84.00\%}$
Visuomotor Student Non-Hierarchical $0.0\%$ $0.0\%$ $0.0\%$ $0.0\%$ $0.0\%$ $0.0\%$ $0.0\%$
Floating Base $42.45\%$ $0.0\%$ $37.36\%$ $8.33\%$ $41.41\%$ $43.57\%$ $3.70\%$
VBC (Ours) $\mathbf{55.40}\%$ $\mathbf{28.57}\%$ $\mathbf{80.00}\%$ $\mathbf{56.57}\%$ $\mathbf{68.01}\%$ $\mathbf{58.96\%}$ $\mathbf{53.33\%}$
@@
Figure 11: Average success times w.r.t. the trial times of picking up 14 objects at different heights, tested in the real world. We allow 5 continuing trial times (including 4 continuous retries, x-axis) during one experiment, i.e., one reset of picking up one specific object. A failure is recorded if the robot attempts more than 5 trails in one reset or if the object falls from a higher surface. For each object and each height, such experiments are conducted 5 times, and therefore we calculate the average success time (y-axis) across these trials, representing the mean number of success times across five trials. It is obvious that the emergent retrying behavior improves the performance.
Evaluation principles and baselines. In the simulation, when the environment is reset and a new trail begins, both the robot and the objects are initialized with random positions and poses, and it’s considered a success if an object is picked up and a failure when the object falls down the table or is not successful in 150 high-level steps. During experiments, we compare both the privileged teacher policies of VBC against baselines, and their visuomotor student policies over $34$ kinds of different objects: 1) VBC w.o. shape feature: a visuomotor policy that is trained the same as VBC, but does not access the pre-trained object shape features during the teacher training stage. This aims to show the necessity of the object shape features during teacher training. We only compare the teacher’s performance as the only difference comes from the teacher’s side. 2) Floating base: a floating base policy with a perfect low-level navigation ability but without whole-body behavior; in other words, the robot can always follow the given velocity commands and body height commands ranging from 0.4 to 0.55 meters (we set 0.4 meters as the lower bound as the default controller in real-world only allows 0.47 meters in minimum). This is to show the indispensability of the whole-body behavior. We compare both the teacher and its corresponding student policy using the same policy network structure. 3) Non-hierarchical: a unified policy trained with low-level policy and visuomotor high-level policy jointly in an end-to-end style. This policy takes the observation of our low-level and high-level policies together. It outputs all the target positions of 12 robot joint angles and the target pose of the gripper. We failed in training such a policy, indicating the effectiveness and necessity of the training pipeline of VBC. We design these baselines to provide a valuable analysis of the design of each module and part.
Figure 12: Visualization of robot behavior and corresponding commands during the test time. Particularly, we plot the two kinds of high-level commands given by the high-level policy, as described in Section III-C, i.e., the velocity commands (including the quadruped linear velocity and yaw velocity); along with the gripper command (close or open).
Picking up different objects. We test the picking-up success rate on each distinct object for more than $300$ episodes and collect the results over every category. During test time, all object heights are randomly set from $0.0$m to $0.6$m for each trial. As is shown in Tab. I, VBC achieves the best performance on 4/7 categories of objects. It is also interesting that with 3D features, VBC becomes worse at regular objects (compared to VBC w.o. Shape Feature), but is much better on irregular objects with complicated shapes, i.e., long box, cup, bowl and drill. One possible reason is that, without pre-trained features, the policy targets the object’s center for grasping, which is effective enough on simpler objects yet inapplicable for objects with more complicated shapes. One may notice that the student policy is much worse than the teacher policy, this is because when the robot touches but fails to grasp on the first try, the object will be randomly laid so that its pose will be much harder to grasp within only 150 steps using only visual inputs. In addition, the performance of the student policy of VBC behaves the worst on long boxes. This is because without shape features, it is much harder to determine the pose of the object, and some of the initial poses of long boxes are extremely hard to grasp for the gripper we used at some height (also as shown in Fig. 10).
Picking up over various heights. We test the mean success rate on each object category on a fixed height, where we also list the object with the maximum/minimum success rate on each height. We compare VBC and the floating base baseline on 7 different heights, and illustrate the results in Fig. 10. From the significant improvement of VBC on almost every height, compared to the floating base policy, we highlight the advantage of VBC that is able to achieve flexible whole-body behaviors in achieving high-level tasks.
IV-B Real-World Experiments
Objects In the real world, we choose 14 objects, including four irregular objects, six common objects, and four regular shapes, shown in Fig. 9. It is worth noting that almost all objects are unseen in the training, i.e., there are no identity objects during real-world experiments.
@@ IV-B Real-World Experiments
Evaluation principles and baselines. We deploy the trained visuomotor student policy of VBC as tested in the simulator directly into the real world. In the real world, we test how many times the robot can pick up the assigned object in 5 trials (i.e., the robot is allowed to continuously retry four times when it fails to grasp without resetting), as we find that the robot emerges to have retrying behavior, as will be described below. We choose this as a measurement since we find that the emergent retrying behaviors improve the robustness, reliability, and success rate of finishing a job. We consider it a success when the object is picked more than 0.1m higher than the plane where it is placed and a failure when the robot tries more than five times or the object falls from its surface. Each object is evaluated over three different heights, as stated before, and five resets are permitted for each height. Regarding baselines, we find it hard to build a fair autonomous method for comparison, so we compared VBC with the default Unitree base controller (allowing the heights of the arm base to vary from 0.47m to 0.55m) combined with a similar high-level policy trained on stationary legs, in which the robot has no whole-body behaviors but also retains the retrying. We teleoperated the robot to be closed enough to the object before the policy started to grasp. During the test of the real-world experiments, we put the objects at a distance that the cameras mounted on the robot can observe, thus they can be manually annotated using our developed TrackingSAM tool. After the annotation, the robot walks by and picks up the objects in a fully autonomous way.
@@
Emergent retrying behaviors. It is worth noting that during simulation training, the robot emerges to learn a retrying behavior. In other words, when the robot fails to grasp an object, it will automatically retry to grasp without any human intervention. This shows an advantage and a clear difference compared to model-based methods. In our observation, as long as the visual mask is on-track/not-lost, the success rates benefit from 1) the change of the object pose in the early tries; and 2) the movement of the robot body.
@@
Qualititive analysis. We want to analyze whether the commands sent by the high-level policy are reasonable to call the low-level policy to achieve the task. To this end, we visualize the velocity commands, the gripper command, along with their corresponding robot behavior during the testing time in Figure 12. It appears that there is less gap when deploying our policy trained in the simulator into the real world, which is able to provide the correct commands that show clear phases.
V Conclusions, Limitations and Future Works
In this paper, we proposed a fully autonomous mobile manipulation system based on quadruped robots and a hierarchical training pipeline, Visual Whole-Body Control (VBC). VBC consists of a low-level goal-reaching policy and a high-level task-planning policy, which is trained through reinforcement learning and imitation learning. Although we showed great success in picking various objects, we found several failure cases due to the limitations of the system: 1) Compounding error. Our pipeline is hierarchical and requires each module to be coherent and precise to make the robot work well. 2) In-precise depth estimation. The RGBD camera does not reflect a perfect depth estimation, especially on reflective objects. 3) Bad gripper. The gripper of the Unitree Z1 Arm is a beak-like gripper instead of a parallel one, which tends to push objects away and is hard for accurate manipulation. 4) Unstable vision inputs. The tracking SAM that is used to provide the mask of annotated objects, may lose track due to camera deviation and occlusion, or be confused when the selected objects have a similar color to the other things. This is the most common cause leading to failure in our real-world experiments. In our current pipeline, before the robot autonomously picks up objects, we must manually specify the target object in both cameras mounted on the robot. Thanks to the generalizability of trackingSAM, we find that as long as the object is visible to one camera (either the head or the wrist camera), the annotation mask can be propagated to another view for automated picking, and free the robot from being specified in the between of the range of the two cameras, which improves the usability of our system. Since the requirements of annotation and object visibility are totally decoupled from our sim-to-real framework, in the future we can further improve and replace these modules with automated annotation algorithms and a search/exploration module to further improve the automation of the whole system. Future works should also consider upgrading the hardware, like replacing a better gripper and RGBD cameras that provide better depth estimation; and simplify the working pipeline.

### v5

@@ 3.2.2 Visuomotor Student Policy
We find robust and adaptive picking-up behaviors emerge from the privileged teacher policy training, with inaccessible object poses and shape features in reality. To deploy our policy in the real world, we must use easily accessible real-world observations like images.
@@
Observations and actions. Although camera images are easily accessible both in simulation and the real world, there is a huge gap for colored images between them. As the main difference and noise come from the surrounding environment in simulated and real images, we turn to segmented depth images to train the visuomotor policy. Formally, the observation of our vision policy is $\mathbf{o}_{t}=[\mathbf{o}^{\text{image}}_{t},\mathbf{s}_{t}^{\text{proprio}},\mathbf{a}_{t-1}]$, where $\mathbf{o}^{\text{image}}_{t}$ consists of object segmentation masks and segmented depth images. $\mathbf{s}_{\text{proprio}}$ and $\mathbf{a}_{t-1}]$ are proprioceptive observation and last action defined above. The output actions are the same as those of the teacher policy.
@@
Imitation learning for student distillation. To deploy our policy in the real world, we distill our state-based high-level planning policy into a vision-based policy using DAgger, which warms up the training by teacher sampling the initial transition data. Then we sample transitions using the student policy and request correction from the teacher. We find this is much more efficient and effective than directly learning a visuomotor student policy using RL.
@@
Randomization and augmentation. To bridge the sim-and-real gap, we slightly randomize the position and rotation of both cameras to handle slight offsets in real. Besides, we clip the depth image to a minimum of 0.2m to avoid low-quality images when the distance is too close to a real camera, after which we normalize the depth value. We adopt augmentations like RandomErasing, GaussianBlur, GaussianNoise, and RandomRotation for the depth images. Additional details regarding the high-level policy training including rewards can be found in Appendix C.2.
3.3 Real World Deployment
Robot Platform The robot platform used in this work comprises a Unitree B1 quadruped robot with a Unitree Z1 robot arm mounted on its back. Our robot equips two cameras, one on the head and another close to the gripper. We use the onboard computer of B1 to control the low-level policy, and an additional onboard computer to receive visual inputs and control the high-level policy. The Unitree B1 quadruped has 12 actuatable DoFs, the arm has 6 DoFs, and the gripper has 1 DoF, making 19 DoFs in total. An overview of the hardware setup is shown in Fig. 4.
@@ 3.3 Real World Deployment
Vision masks. Note that our system works with little human intervention, i.e., we need to annotate and segment objects from environment backgrounds. To this end, we take use of TrackingSAM [24] (details in Appendix J), a tool that combines a video object tracking model [25] with interactive segmentation interface [26]. The implementation details and an example for illustration are given in Sec. J and Fig. 14. During real-world experiments, we annotated the objects from one of the RGBD cameras at the beginning of each reset trial. TrackingSAM will keep tracking the segmented object and providing the mask in real time ($\sim$10 fps on the robot). The depth images obtained from the RGBD cameras are pre-processed by hole-filling and clipping. Empirically, as long as the object is visible to one camera (either the head or the wrist camera), the annotation mask can be propagated to both views to initialize an automated pickup.
4 Experiments
@@
We conduct a set of experiments on pick-up tasks to compare VBC with baselines both in simulation and the real world, showing the effectiveness of training a loco-manipulation policy with VBC, and the generalizability over varying objects with different heights.
Experiment setups. We utilize IsaacGym [27] for parallel training. In the real-world experiment, we initialize objects in a random location and orientation in front of the robot, making sure it is visible in at least one camera (around 0.5$\sim$1.5m). We test three heights by putting the objects on the ground, on a box, and on a table, corresponding to the heights of 0.0m, $\sim$0.3m, $\sim$0.5m.
Objects. We use $33$ YCB objects [28] in simulation. According to the object shapes, we roughly divided them into $7$ representative categories: ball, long box, square box, bottle, cup, bowl and drill. As for real-world experiments, we choose 14 objects, including 4 irregular objects, 6 common objects, and 4 regular shapes. It is worth noting that almost all objects are unseen in the simulation. See Appendix H for illustrated details.
4.1 Simulation Results
Figure 6: Success rates of different student policies at different heights in simulation. We test each distinct object for more than 100 episodes. The dots/crosses represent the mean performance, and the upper/lower bound is the max/min success rate over 7 categories, with the category name noted. The non-hierarchical baseline is not included for not reaching meaningful results.
Evaluation principles and baselines. In the simulation, we consider a successful grasp if the object is picked up and a failure when the object falls down the table or is not successful in 150 high-level steps ($\sim 18-24s$). We compare both the privileged teacher policies of VBC against baselines and their visuomotor student policies to provide an analysis of each module of VBC: 1) VBC w.o. shape feature: a visuomotor policy that is trained the same as VBC, but does not access the pre-trained object shape features during the teacher training stage. 2) Floating base: a floating base policy with a perfect low-level navigation ability but without arm-leg coordination; in other words, the robot can always follow the given velocity commands and body height commands ranging from 0.4 to 0.55 meters (we set 0.4 meters as the lower bound as the default controller in the real-world only allows 0.47 meters in minimum). We compare both the teacher and its corresponding student policy using the same policy network structure. 3) Non-hierarchical: a unified policy trained with low-level policy and visuomotor high-level policy jointly in an end-to-end style. This policy takes the observation of our low-level and high-level policies together and outputs all the target positions of 12 robot joint angles and the target pose of the gripper. We failed in training such a policy, indicating the effectiveness and necessity of VBC’s training pipeline.
Table 1: Success rates of VBC compared to baseline method on seven categories of objects, tested in the simulator. Each distinct object is tested for more than 300 episodes over 3 random seeds. The robot starts from a 1.0-meter distance from the table. Every episode ends when the robot successfully picks up the object, the object falls from the table, or the robot fails until 150 steps. With the 3D shape feature, the algorithm is better at handling irregular objects.
@@ 4.1 Simulation Results
Privleged Teacher Floating Base $0.63\pm 0.01$ $0.20\pm 0.03$ $0.73\pm 0.01$ $0.36\pm 0.02$ $0.71\pm 0.02$ $0.71\pm 0.01$ $0.07\pm 0.02$
VBC w.o. Shape Feature $\mathbf{0.96\pm 0.01}$ $0.74\pm 0.03$ $\mathbf{0.96\pm 0.00}$ $\mathbf{0.82\pm 0.01}$ $0.84\pm 0.01$ $0.55\pm 0.01$ $0.74\pm 0.00$
VBC (Ours) $0.89\pm 0.01$ $\mathbf{0.93\pm 0.00}$ $81.87\pm 0.02$ $0.76\pm 0.01$ $\mathbf{0.89\pm 0.02}$ $\mathbf{0.78\pm 0.03}$ $\mathbf{0.84\pm 0.00}$
Visuomotor Student Non-Hierarchical $0.0$ $0.0$ $0.0$ $0.0$ $0.0$ $0.0$ $0.0$
Floating Base $0.39\pm 0.07$ $0.06\pm 0.02$ $0.42\pm 0.02$ $0.15\pm 0.03$ $0.38\pm 0.05$ $0.44\pm 0.01$ $0.03\pm 0.01$
VBC (Ours) $\mathbf{0.67\pm 0.07}$ $\mathbf{0.42\pm 0.12}$ $\mathbf{0.74\pm 0.09}$ $\mathbf{0.55\pm 0.01}$ $\mathbf{0.74\pm 0.06}$ $\mathbf{0.52\pm 0.02}$ $\mathbf{0.47\pm 0.11}$
Picking up different objects. We test the picking-up success rate on each distinct object for more than $300$ episodes and collect the results over every category. As is shown in Tab. 1, VBC achieves the best performance on 4/7 categories of objects. It is also interesting that with 3D features, VBC becomes worse than VBC w.o. Shape Feature at regular objects, but is much better on irregular objects with complicated shapes, i.e., long box, cup, bowl, and drill. One possible reason is that, without pre-trained features, the policy targets the object’s center for grasping, which is effective enough on simpler objects yet inapplicable for objects with more complicated shapes. One may notice that the student policy is much worse than the teacher policy, since when the robot touches but fails to grasp on the first try, the object may be randomly laid so that its pose will be much harder to grasp within only 150 steps using only visual inputs. In addition, the performance of the student policy of VBC behaves the worst on long boxes. This is because without shape features, it is much harder to determine the pose of the object, and some initial poses of long boxes are extremely hard to grasp for the gripper we used at some height (also as shown in Fig. 6).
Picking up over various heights. We test the mean success rate on each object category on a fixed height, where we also list the object with the maximum/minimum success rate on each height. We compare VBC and the floating base baseline on 7 different heights, and illustrate the results in Fig. 6. From the significant improvement of VBC on almost every height, compared to the floating base policy, we highlight the advantage of VBC in achieving flexible whole-body behaviors.
4.2 Real-World Experiments
@@
Figure 7: Average success times w.r.t. the attempts of picking up 14 objects at different heights, tested in the real world. We allow 5 continuing attempts (including 4 continuous retries, x-axis) during one configuration, i.e., one reset of picking up one specific object at a specific height. A failure is recorded if the robot attempts more than 5 attempts in one reset or if the object falls from its surface. We test 5 times for each object and each height, and calculate the average success time (y-axis) across these trials. The emergent retrying behavior improves the performance.
@@
Evaluation principles and baselines. We deploy the trained visuomotor student policy of VBC directly into the real world. Particularly, we test how many times the robot can pick up the assigned object in five attempts without resetting (i.e., the robot is allowed to continuously retry four attempts when it fails to grasp), since we find that the emergent retrying behaviors improve the robustness, reliability, and success rate of finishing a job. We consider it a success when the object is picked more than 0.1m higher than the placed plane and a failure when the robot tries more than five attempts or the object falls from its surface. Each object is evaluated over three different heights and we do five resets for each configuration. Regarding baselines, as it is hard to build a fair autonomous method for comparison, we compared VBC with teleoperating the default Unitree base controller (allowing the heights of the arm base to vary from 0.47m to 0.55m) combined with a similar high-level policy trained on stationary legs, in which the robot has no arm-leg coordination but also retains the retrying. For this baseline, We teleoperated the robot to be close enough to the object before grasping it. During all tests, we put objects at a distance that one mounted camera can observe, thus they can be easily manually annotated using the TrackingSAM tool. After the annotation, the robot walks towards and picks up the objects in a fully autonomous way. In experiments, we observe that the policy is capable of recovering from short-term tracking loss, which is not uncommon due to the robot turning or tilting.
@@
Emergent retrying behaviors. The robot emerges to learn a retrying behavior in simulation training. In other words, when the robot fails to grasp an object, it automatically retries to grasp without human intervention. This shows an advantage and a clear difference compared to model-based methods. In our observation, as long as the visual mask is on-track/not-lost, the success rates benefit from 1) the change of the object pose in the early tries; and 2) the movement of the robot body.
@@
Figure 8: Visualization of robot behavior and corresponding commands during the test time. We plot the velocity commands (including the quadruped linear velocity and yaw velocity); along with the gripper command (close or open).
Qualititive analysis. We analyze whether the commands sent by the high-level policy are reasonable to call the low-level policy to achieve the task. To this end, we visualize the velocity commands, the gripper command, along with their corresponding robot behavior during the testing time in Fig. 8. Our system has small gaps when deploying our policy trained in the simulator into the real world, which is able to provide the correct commands that show clear phases.
5 Conclusions and Limitations
We proposed Visual Whole-Body Control (VBC), a fully autonomous loco-manipulation system on quadruped robots with a hierarchical training pipeline. VBC trains a low-level goal-reaching policy and a high-level task-planning policy through reinforcement learning and imitation learning.
Limitations. Although we showed great success in picking various objects, there are several limitations regarding the system and the tasks that can be of interest in future work. A detailed analysis is in Appendix I and we outline some reasons: (1) visual tracking loss; (2) inadequate gripper design; (3) in-precise depth estimation; (4) compounding error.